John is twice as old as his brother. Four years ago, he was four times as old as his brother. find their present ages.
Answer by rapaljer(4671)

Let x= brother's age now
2x = John's age now
x-4 = brother's age 4 years ago
2x-4 = John's age 4 years ago

The EQUATION:
John's age 4 years ago was four times the brother's age 4 years ago.
2x-4=4(x-4)
2x-4=4x-16

Subtract 2x from each side:
2x-2x-4=4x-2x-16
-4= 2x-16

Add + 16 to each side:
16-4=2x-16+16
12=2x
x=6

Check:
Brother = 6 now
John = 12 now

Four years ago, John was 8 and Brother was 2, which is 4 times as old.

Dr. Robert J. Rapalje, Retired